OpenStudy (anonymous):

Assume Pr[A U B]=.75 and Pr[A]=.3 What is Pr[B] if A and B are independent?

OpenStudy (anonymous):

Pr(A U B) = Pr(A) + Pr(B) - Pr(A intersect B) A and B are independent, so what does the above equation simplify to? From there, you should have your answer.

OpenStudy (anonymous):

Is A intersection B just Pr[A]?

OpenStudy (anonymous):

Close, but not quite. A intersect B are the elements that are in A that are also in B. But, the two sets are independent, which means there is no element which is in both. I can draw a picture if you'd like.

OpenStudy (anonymous):

If they are disjoint, then Pr(A intersect B)=0 and Pr(B)=Pr(A U B)-Pr(A)=.45, but it keeps saying that is wrong and that is only if they are disjoint, not independent

OpenStudy (anonymous):

You're right, sorry. I'm mistaken. I only took one probability course as an undergraduate. Pr(A intersect B) = Pr(A)*Pr(B) if A and B are independent.

OpenStudy (anonymous):

I'm not sure how I would find B from what I have though

OpenStudy (anonymous):

P(A U B) = P(A) + P(B) - P(A intersect B) = P(A) + P(B) - P(A)*P(B) So, 0.75 = 0.3 + P(B) - 0.3*P(B) => 0.45 = 0.7P(B) P(B) = 0.6428 (rounded)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

You're welcome. Sorry for the confusion.

OpenStudy (anonymous):

No worries!